# Scaling of energy delivered through an electrostatic discharge to a small series load

### v1

@@ Abstract
We study the energy delivered through a small-resistance series “victim” load during electrostatic discharge events in air. For gap lengths over 1 mm, the fraction of the stored energy delivered is mostly gap-length independent, with a slight decrease at larger gaps due to electrode geometry. The energy to the victim scales linearly with circuit capacitance and victim load resistance but does not strongly dependent on circuit inductance. This scaling leads to a simple approach to predicting the maximum energy that will be delivered to a series resistance for the case where the victim load resistance is lower than the spark resistance.
@@ 1 Introduction
Advances in the design of electronic devices are often accompanied by an increased sensitivity to electrostatic discharge (ESD). ESD damage to a device can cause direct failure due to physical damage and/or indirect failure due to the presence of strong electric fields or transient electric fields [1, 2]. Wearable electronics have increased in popularity, with particular interest in the risk of ESD [3]. Although voltage or current thresholds are often used as a measure of risk to electronic devices, for short ESD pulses, the energy delivered from an ESD event is often a good measure of risk to the device [4].
@@
In other settings, the risk extends beyond equipment damage: ESD can initiate combustion of energetic materials, resulting in serious consequences [5, 6, 7, 8, 9, 10, 11, 12]. The “ESD sensitivity”, the necessary energy delivered to the energetic material to cause it to ignite 50% of the time, is often used as a measure of the overall sensitivity of the material. This has shown good agreement for different experiments with many materials [10, 13]. Still, the sensitivity to ESD can also depend on several parameters, such as chemical composition, granularity and grain shape, mechanical properties, and temperature and moisture content [9].
@@
Various studies have been performed to better understand ESD events in a variety of experimental configurations, such as a spherical-plane geometry [14, 15, 16, 17, 18]. Current delivery from charged humans is of particular interest [19, 20]. Some research has been conducted in transient regimes, for example, to study the effect of approach speed to a grounded electrode on peak currents and rise times [21, 22, 23] or for fast voltage rise times [24].
@@
In this paper, we consider a scenario in which two conducting electrodes are brought to breakdown in a quasi-static fashion, causing an ESD event in the air between these electrodes. The “spark channel” formed between the electrodes is considered to be in series with a “victim” load. The victim could be a sensitive electronic component or a combustible material. From the perspective of energy dynamics, energy is expended ionizing and exciting the gas between the electrodes, forming a plasma channel, Ohmic heating of the plasma channel, hydrodynamic expansion often forming a shock wave, and radiation of electromagnetic waves. The remaining energy is delivered to the victim load. Understanding how energy is partitioned between the spark channel and victim load may help guide safety requirements set around sensitive devices and materials.
@@
We employ a well-characterized circuit to understand the energy transfer to a victim load under different circuit parameters. Several electrode geometries and circuit capacitances, inductances, and gap lengths were tested. This allows us to gain insight into how the energy transfer scales with these parameters and to identify different physical regimes that occur for ESD events of varying stored energy. We measure current and voltage profiles over the entire discharge current pulse and use this to determine the energy delivered to the spark and victim. Building on our previous theoretical analysis [25], we find that the spark resistance model initially put forth by Rompe and Weizel [26] performs quite well in predicting the scaling of the energy transfer.
@@ 2 Material and Methods
We constructed two different setups: an open-air system (OAS) and a symmetric vacuum chamber (SVC). The SVC spark gap electrode configuration can accommodate electrodes of different shapes: 2.00 cm diameter graphite spheres, 1.27 cm diameter brass spheres, a chrome-plated steel needle (0.787 mm tip diameter), or a 1.59 cm diameter flat steel disk. The capacitances of the electrode systems were estimated to be below 10 pF. The OAS system spark gap is modified from a commercial high-voltage spark gap (Ross Engineering SG-40-H), Figure 2(a). This spark gap has two 3.75 cm diameter graphite spherical electrodes separated by a variable gap length. We inferred the inductance of the OAS system to be approximately 1.1 $\mu$H from the frequency of the ring-down of the discharge current. This relatively large inductance results from the size of the current loop created by the return wire. Finite element simulation (using Comsol Multiphysics) of the discharge circuit gave an inductance of 0.98 $\mu$H, in reasonable agreement with the measured inductance.
Figure 2: Graphical depiction of the discharge circuit elements $C_{x}$, CVR ($R_{v}$) and spark gaps of the (a) open-air system (OAS) and the (b) symmetric vacuum chamber (SVC). For the SVC, the image is a cross-sectional view of the chamber to visualize the circuit components
The SVC system is an in-house built vacuum chamber designed and manufactured for these ESD experiments. A cross-sectional view is shown in Figure 2(b). The outside of the chamber was designed to provide a much more symmetric return to the ground to reduce the system inductance. During the design process, the finite element models predicted an inductance of 142 nH; the measured system inductance was 120 nH. Like the OAS, this system allows for a variable gap length and the ability to change electrodes. Since the chamber is sealed (brass and viewing windows), there is the capability to control the gas composition. For the data presented here, all gas compositions are in air at standard pressures for Golden, Colorado (roughly 630 torr or 0.83 atm).
@@
Several circuit parameters were varied to investigate the effects on energy transferred to the victim load through an ESD: external capacitance ($C_{x}$), electrode geometry, and inductance of the circuit (taking advantage of the different inductances of the OAS and SVC systems). On the SVC system, the external capacitance was varied from 100 to 700 pF to test how the discharge energy changes the fraction of energy delivered through the discharge channel. The SVC was also used to examine the effects of varying electrode geometry. The different geometries used were sphere-sphere, plane-sphere, and needle-plane. For asymmetric electrode geometries, the anode-cathode polarity was also switched.
@@ 2.1 Data Processing
Figure 3: SVC example voltage and current traces with $C_{x}=100$ pF, 0.0983 $\Omega$ CVR, 1.27 cm diameter brass spherical electrodes and Cal Test voltage probe. The legend in (a) applies to both plots.
@@
$Q_{T}$ $=\sum_{t=0}^{t}I(t)*\Delta t$ $\mathcal{E}_{T}=\sum_{t=0}^{t}I(t)V(t)*\Delta t$ (2)
Here $V(t)$ is the signal from the HVP and $\Delta t$ is the time step. $Q_{0}$ and $\mathcal{E}_{0}$ are the initial charge and stored energy, respectively, corresponding to the measured breakdown voltage. $Q_{T}$ and $\mathcal{E}_{T}$ are the charge and energy obtained by integrating the current and power delivered through the spark and victim load throughout the discharge event.
Figure 4 shows the average current and integrated charge traces for 20 discharge events using the SVC with a 100 pF $C_{x}$ and 0.0983 $\Omega$ CVR at a gap length of $h=4.2$ mm. The dynamics show that the charge transfer peaks after the first half of the current cycle.
Figure 4: The average current (dashed line) and integrated charge (solid line) for 20 individual discharge events, using the SVC with a 100 pF $C_{x}$ and 0.0983 $\Omega$ CVR at a 4.2 mm gap. The stored charge $Q_{0}$ is indicated by the horizontal line and marked on the Charge axis.
@@
The importance of this phase measurement is illustrated in Figure 5, which shows the power through the discharge circuit and the integrated energy for 20 discharge events using the SVC with 100 pF $C_{x}$ and 0.0983 $\Omega$ CVR at a 3.0 mm gap. Figure 5(b) confirms that the measured phase delay of $-2.2$ ns yields the correct integrated energy. It also shows that the majority of the energy is transferred within the first half cycle of the current pulse.
Figure 5: For 20 individual discharge events, shown here is the (a) voltage (solid lines) and current trace (dashed line), and the (b) dissipated energy through the circuit, using the SVC with $C_{x}=100$ pF and 0.0983 $\Omega$ CVR at a 3.0 mm gap. The analysis is shown for three different cable delays, emphasizing the importance of the time shift between the voltage and current signals. The shaded areas around the curves show the spread in signal, and the horizontal line in (b) indicates the stored energy. The legend in (a) applies to both plots.
The stored and dissipated charge and energy as a function of the gap length are shown in Figure 6, for various external capacitors in the SVC. In the figure, the stored charge and energy ($Q_{0}$ and $\mathcal{E}_{0}$) are the solid lines, and the dissipated charge and energy ($Q_{T}$ and $\mathcal{E}_{T}$) are the points. The conservation of charge and energy was verified for all data presented here.
Figure 6: Comparison between the stored (solid line) and integrated (points) charge and energy for different external capacitors in the SVC with a 0.25 $\Omega$ CVR. The legend in (b) applies to both plots.
In these experiments, we used the current-viewing resistor (CVR) as our victim load. We used several different CVR resistances (0.005140 to 0.2505 Ohms) and found that these small values did not significantly affect the discharge current. Therefore, in the following data, we normalized the energy delivered to the victim load to a victim resistance of 0.1 Ohm.
The victim energy, $\mathcal{E}_{v}$, is calculated using the CVR traces. The power through the victim load is $P=I_{v}^{2}R_{v}$, and the total energy delivered to the victim is obtained by integrating the power, see Eq. (8).
@@ 3 Theory
To interpret the measurements of energy transferred to the victim load, we will make use of a simple model of nonlinear spark resistance originally presented by Rompe and Weizel [26]. The Rompe-Weizel (RW) model, along with several other circuit-based models, are frequently used as the starting point for ESD models today [25]. The RW model is based on the assumption that there is a certain energy cost per electron created by the breakdown process ($U_{\text{eff}}$) and that the mobility ($\mu$) of the electrons is constant.
In this simple model, we assume that the electric field is uniform within the gap so that the threshold voltage for breakdown, $V_{\text{th}}$, is linearly proportional to the gap length, $V_{\text{th}}=E_{\text{th}}h$. Here, the electric field strength for breakdown, $E_{\text{th}}$, can be obtained from measured data found in Paschen curves. The energy required to generate electrons in a low-temperature plasma is, in practice, much higher than the actual ionization energy of the background gas. For the purposes of this work, $U_{\text{eff}}$ is considered an empirical parameter, as it includes a variety of energy sinks (ionization and dissociation energy, particle heating and excitation). Joule heating, $\vec{J}\cdot\vec{E}$, transfers energy from the electromagnetic field to the plasma, which is assumed to balance with the energy expended to create the electrons:
@@
$R_{S}(t)=\left(\frac{2\,a_{R}}{h^{2}}\int_{0}^{t}I(t^{\prime})^{2}\mathrm{d}t\right)^{-1/2},$ (4)
@@
For this research, we are interested in the energy delivered to the victim load. In the case of uniform electric field, the initial stored energy ideally increases quadratically with $h$:
@@
In this limit, the final resistance is independent of the spark gap separation, $h$.
The energy delivered to the victim load, $\mathcal{E}_{v}$, is obtained by integrating the power:
$\mathcal{E}_{v}=R_{v}\int_{0}^{\infty}I^{2}\mathrm{d}t.$ (8)
@@
$\mathcal{E}_{v}=R_{v}\frac{h^{2}}{2\,a_{R}\,R_{\text{SF}}^{2}}.$ (9)
In the simple limit where the stored energy follows Eq. (5), we obtain:
$\mathcal{E}_{v}=\frac{R_{v}}{R_{\text{SF}}^{2}}\frac{\mathcal{E}_{0}}{a_{R}\,C\,E_{\text{th}}^{2}}.$ (10)
In the limit of small victim load resistance ($R_{v}$), most of the energy is delivered to the spark, and $R_{\text{SF}}=R_{\text{SFmin}}$, so $\mathcal{E}_{v}=\frac{1}{2}\frac{R_{v}}{R_{\text{SFmin}}}\mathcal{E}_{0}$ and the fraction of stored energy delivered to the victim load is:
$\eta_{v}=\frac{\mathcal{E}_{v}}{\mathcal{E}_{0}}=\frac{1}{2}\frac{R_{v}}{R_{\text{SFmin}}}=\frac{1}{4}a_{R}\,C\,R_{v}\,E_{\text{th}}^{2}.$ (11)
When looking at the fractional energy, we can move all circuit parameters to one side to obtain
$\overline{\eta}_{v}=\frac{\eta_{v}}{CR_{v}}=\frac{a_{R}E_{\text{th}}^{2}}{4}.$ (12)
This expression shows that if the fraction of stored energy is scaled by $R_{v}$ and $C$, the result will be independent of the gap length. We will test this result in Section 4 below.

### v2

@@ Abstract
We study the energy delivered through a small-resistance series “victim” load during electrostatic discharge events in air. For gap lengths over 1 mm, the fraction of the stored energy delivered is mostly gap-length independent, with a slight decrease at larger gaps due to electrode geometry. The energy to the victim scales linearly with circuit capacitance and victim load resistance but is not strongly dependent on circuit inductance. This scaling leads to a simple approach to predicting the maximum energy that will be delivered to a series resistance for the case where the victim load resistance is lower than the spark resistance.
@@ 1 Introduction
Advances in the design of electronic devices are often accompanied by an increased sensitivity to electrostatic discharge (ESD). The effects of ESD on a device can be direct, causing physical damage and immediate failure, or indirect, leading to delayed failure from strong or transient electric fields [Smallwood2020, Greason1984]. The recent popularity of wearable electronics has drawn particular attention to their ESD risk [Zhou2018]. Although voltage or current thresholds are often used as a measure of risk to electronic devices, the energy delivered by a short ESD pulse is often a good measure of risk to the device [Paasi2005].
@@
In other settings, the risk extends beyond equipment damage: ESD can initiate combustion of energetic materials, resulting in serious consequences [Lovstrand1981, Larson1989, Guoxiang1982, Rizvi1992, Hearn2001, Talawar2006, Ohsawa2011, Weir2013]. The “ESD sensitivity”, the energy required to ignite energetic material 50% of the time, is often used as a measure of the overall sensitivity of the material. This metric has shown good agreement for different experiments with many materials [Larson1989, Skinner1998]. Still, the sensitivity to ESD can also depend on several parameters, such as chemical composition, granularity and grain shape, mechanical properties, and temperature and moisture content [Talawar2006].
@@
Various studies have been performed to better understand ESD events in a variety of experimental configurations, such as a spherical-planar geometry [Zaridze1996, Bach1981, Hearn2005, Mori2007, Chubb1982]. Specific attention has been given to the current delivered from charged human bodies [Katsivelis2015, Taka2009] as well as to transient phenomena, such as the effects of approach speed [Yoshida2012, Daout1987, Tomita2015] or fast voltage rise times [Parkevich2019] on peak current and peak power.
The efficiency of energy transfer in a spark discharge, defined as the amount of stored energy converted into spark energy, has been a key area of study, particularly concerning gas ignition [Zhong2015, Peng2013, Eckhoff2010, Randeberg2006]. Zhong et al. [Zhong2015] investigated the influence of different trigger methods and inductance loads (0.024 mH and 1.454 mH) on discharge efficiency and time. Their methodology involved collecting high-voltage and current probe traces to determine spark energy. They employed two primary trigger methods: a pneumatic piston that rapidly moved a grounded electrode toward a fixed high-voltage electrode, which resulted in a discharge at an unknown gap length, and a high-voltage relay that applied an overvoltage to a 6 mm gap. Within this research, only transient ESD events were studied.
In an experimental investigation of spark discharge energy, Peng et al. [Peng2013] determined the energy delivered to the spark using an integration method of measured voltage and current in a needle-to-plate electrode configuration. Their study allowed for double current pulses, increasing the discharge efficiency to nearly 100%.
[…]
@@
In this paper, we consider a scenario in which two conducting electrodes are brought to breakdown in a quasi-static fashion, causing an ESD event in the air between these electrodes. The “spark channel” formed between the electrodes is treated as a resistive element in series with a “victim” load. The victim could be a sensitive electronic component or a combustible material. From the perspective of energy dynamics, energy is expended ionizing and exciting the gas between the electrodes, forming a plasma channel, Ohmic heating of the plasma channel, hydrodynamic expansion often forming a shock wave, and radiation of electromagnetic waves. The remaining energy is delivered to the victim load. In the classic Rompe-Weizel (RW) model, the spark resistance is assumed to change only due to the increase in electron density. All of the remaining physics is gathered into an empirical parameter ($a_{R}$) that corresponds to the ratio of the electron mobility to the net energy required to produce new electrons. We show that this simple model performs surprisingly well, in spite of the many mechanisms for sinking energy into the spark channel. Understanding the energy delivery to a victim load is critical for a wide range of engineering applications. This research is particularly valuable for guiding safety requirements for sensitive devices and materials with small resistance, such as those used in explosive bridge wires as well as common devices that rely on low-resistance components, such as the shunt resistors in multimeters or the circuits within inductive cooktops. The predictions based on the RW model of how energy is partitioned between the spark channel and victim load may help guide safety requirements set around sensitive devices and materials.
@@
In the work presented here, we employ a well-characterized circuit to understand the energy transfer to a victim load under different circuit parameters (for further information see [SchramaCAM2023IoED]). Several electrode geometries and circuit capacitances, inductances, and gap lengths were tested. This allows us to gain insight into how the energy transfer scales with these parameters and to identify different physical regimes that occur for ESD events of varying stored energy. We measure current and voltage profiles over the entire discharge current pulse and use this to determine the energy delivered to the spark and victim. Building on our previous theoretical analysis [Rose2020], we find that the spark resistance model initially put forth by Rompe and Weizel [Rompe1944] performs quite well in predicting the scaling of the energy transfer.
@@ 2 Material and Methods
[…]
Figure 2: Graphical depiction of the discharge circuit elements $C_{x}$, CVR ($R_{\text{v}}$) and spark gaps of the (a) open-air system (OAS) and the (b) symmetric grounded system (SGS). For the SGS, the image is a cross-sectional view of the chamber to visualize the circuit components
The SGS system is an in-house built circuit chamber designed and manufactured for these ESD experiments. A cross-sectional view is shown in Figure 2(b). The outside of the chamber was designed to provide a much more symmetric return to the ground to reduce the system inductance. During the design process, the finite element models predicted an inductance of 142 nH; the measured system inductance was 120 nH. Like the OAS, this system allows for a variable gap length and the ability to change electrodes. Since the chamber is sealed (brass and viewing windows), there is the capability to control the gas composition. For the data presented here, all gas compositions are in air at standard pressures for Golden, Colorado (roughly 630 torr or 0.83 atm at 28 $\pm$ 3% humidity). As a comparison, we obtained very similar results for a dry synthetic air ($\text{N}_{2}-\text{O}_{2}$) mixture.
@@
Several circuit parameters were varied to investigate the effects on energy transferred to the victim load through an ESD: external capacitance ($C_{x}$), electrode geometry, and inductance of the circuit (taking advantage of the different inductances of the OAS and SGS systems). On the SGS system, the external capacitance was varied from 100 to 700 pF to test how the discharge energy changes the fraction of energy delivered through the discharge channel. The SGS was also used to examine the effects of varying electrode geometry. The different geometries used were sphere-sphere, plane-sphere, and needle-plane. For asymmetric electrode geometries, the anode-cathode polarity was also switched.
@@ 2.1 Data Processing
Figure 3: SGS example voltage and current traces with $C_{x}=100$ pF, $R_{\text{v}}=0.0983$ $\Omega$, 1.27 cm diameter brass spherical electrodes and Cal Test voltage probe. The legend in (a) applies to both plots.
Figure 4: The average current (dashed line) and integrated charge (solid line) for 20 individual discharge events, using the SGS with a 100 pF $C_{x}$ and 0.0983 $\Omega$ CVR at a 4.2 mm gap. The stored charge $Q_{0}$ is indicated by the horizontal line and marked on the charge axis.
@@
$Q_{T}$ $=\int_{t=0}^{t}I(t)\,\text{d}t$ $\mathcal{E}_{T}=\int_{t=0}^{t}I(t)V(t)\,\text{d}t$ (2)
Here $V(t)$ is the signal from the HVP. $Q_{0}$ and $\mathcal{E}_{0}$ are the initial charge and stored energy, respectively, corresponding to the measured breakdown voltage. $Q_{T}$ and $\mathcal{E}_{T}$ are the charge and energy obtained by integrating the current and power delivered through the spark and victim load throughout the discharge event.
Figure 4 shows the average current and integrated charge traces for 20 discharge events using the SGS with a 100 pF $C_{x}$ and 0.0983 $\Omega$ $R_{\text{v}}$ at a gap length of $h=4.2$ mm. The dynamics show that the charge transfer peaks after the first half of the current cycle.
@@
The importance of this phase measurement is illustrated in Figure 5, which shows the power through the discharge circuit and the integrated energy for 20 discharge events using the SGS with 100 pF $C_{x}$ and 0.0983 $\Omega$ $R_{\text{v}}$ at a 3.0 mm gap. Figure 5(b) confirms that the measured phase delay of 2.2-2.2 ns yields the correct integrated energy. It also shows that the majority of the energy is transferred within the first half cycle of the current pulse.
Figure 5: For 20 individual discharge events, shown here is the (a) voltage (solid lines) and current trace (dashed line), and the (b) dissipated energy through the circuit, using the SGS with $C_{x}=100$ pF and $R_{v}=0.0983$ $\Omega$ at a 3.0 mm gap. The analysis is shown for three different cable delays, emphasizing the importance of the time shift between the voltage and current signals. The shaded areas around the curves show the spread in signal, and the horizontal line in (b) indicates the stored energy. The legend in (a) applies to both plots.
The stored and dissipated charge and energy as a function of gap length are shown in Figure 6, for various external capacitors in the SGS. In the figure, the stored charge and energy ($Q_{0}$ and $\mathcal{E}_{0}$) are the solid lines, and the dissipated charge and energy ($Q_{T}$ and $\mathcal{E}_{T}$) are the points. The conservation of charge and energy was verified for all data presented here.
For all data points presented, the reported value represents the mean of 10 to 20 unique discharge events, with the error bars indicating the standard deviation. The data point for the 3 mm discharge at 700 pF in Figure 6 exhibited notably larger error bars compared to the other measurements. This increased uncertainty may be attributed to a non-measurable discharge path, potentially caused by compromised electrical contacts within the circuit, such as at the electrodes or ground connections. This alternative path bypassed the CVR, which did not record the full discharge current.
Figure 6: Comparison between the stored (solid line) and integrated (points) charge and energy for different external capacitors in the SGS with a 0.25 $\Omega$ CVR. The legend in (b) applies to both plots. There is one data point with larger error bars, which we attribute to non-ideal conditions such as dust or other dirty connections.
With careful consideration of cable time delays and parasitic circuit inductance, the spark resistance was calculated. We found that for circuits with a small victim load resistance ($R_{\text{v}}=0.0983$ $\Omega$) the spark resistance reached $2.08\pm 0.13$ $\Omega$ with $C_{x}=700$ pF, and 30–50 $\Omega$ with $C_{x}=100$ pF. For the case of this small victim load resistance, $R_{\text{v}}\ll R_{s}$. Further details on these spark resistance measurements, including a broader range of resistance loads, are available in our related work [schrama2025variableRv, SchramaCAM2023IoED].
In these experiments, we used the current-viewing resistor (CVR) as our victim load. We used several different CVR resistances (0.005140 to 0.2505 $\Omega$) and found that these small values did not significantly affect the discharge current. Therefore, in the following data, we normalized the energy delivered to the victim load to a victim resistance of 0.1 $\Omega$.
The victim energy, $\mathcal{E}_{\text{v}}$, is calculated using the CVR traces. The power through the victim load is $P=I_{v}^{2}R_{\text{v}}$, and the total energy delivered to the victim is obtained by integrating the power, see Eq. (8).
@@ 3 Theory
To interpret the measurements of energy transferred to the victim load, we will make use of a simple model of nonlinear spark resistance originally presented by Rompe and Weizel [Rompe1944]. The Rompe-Weizel (RW) model, along with several other circuit-based models, are frequently used as the starting point for ESD models today [Rose2020]. The RW model is based on the assumption that there is a certain energy cost per electron created by the breakdown process ($U_{\text{eff}}$) and that the mobility ($\mu$) of the electrons is constant. The simplicity of models such as the RW model allows for quick calculations. The model does not explicitly consider complicated chemistry or whether energy is delivered to heat, molecular excitation, ionization, radiation, etc. The only assumption of the model is that there is an “effective cost” to generate an electron, which encompasses all possible energy channels. This cost is averaged over all the electrons, making the effective cost constant.
In this simple model, we assume that the electric field is uniform within the gap so that the threshold voltage for breakdown, $V_{\text{th}}$, is linearly proportional to the gap length, $V_{\text{th}}=E_{\text{th}}h$. Here, the electric field strength for breakdown, $E_{\text{th}}$, can be obtained from measured data found in Paschen curves [Paschen1889, Raizer1991]. The energy required to generate electrons in a low-temperature plasma is, in practice, much higher than the actual ionization energy of the background gas. For the purposes of this work, $U_{\text{eff}}$ is considered an empirical parameter, as it includes a variety of energy sinks (ionization and dissociation energy, particle heating and excitation). Joule heating, $\vec{J}\cdot\vec{E}$, transfers energy from the electromagnetic field to the plasma, which is assumed to balance with the energy expended to create the electrons:
@@
$R_{S}(t)=\left(\frac{2\,a_{R}}{h^{2}}\int_{0}^{t}I(t^{\prime})^{2}\mathrm{d}t^{\prime}\right)^{-1/2},$ (4)
@@
For this research, we are interested in the energy delivered to the victim load. In the case of a uniform electric field, the initial stored energy ideally increases quadratically with $h$:
@@
In this limit, the final resistance is independent of the spark gap separation, $h$. For a standard $E_{\text{th}}$ of $30$ kV/cm, $C=700$ pF, and using the listed values for $a_{R}$, $R_{\text{SFmin}}$ is calculated to be in the range of 1.59 to 6.35 $\Omega$. This theoretical minimum spark resistance is therefore greater than the $R_{\text{v}}$ used in these experiments.
The energy delivered to the victim load, $\mathcal{E}_{\text{v}}$, is obtained by integrating the power:
$\mathcal{E}_{\text{v}}=R_{\text{v}}\int_{0}^{\infty}I^{2}\mathrm{d}t.$ (8)
@@
$\mathcal{E}_{\text{v}}=R_{\text{v}}\frac{h^{2}}{2\,a_{R}\,R_{\text{SF}}^{2}}.$ (9)
In the simple limit where the stored energy follows Eq. (5), we obtain
$\mathcal{E}_{\text{v}}=\frac{R_{\text{v}}}{R_{\text{SF}}^{2}}\frac{\mathcal{E}_{0}}{a_{R}\,C\,E_{\text{th}}^{2}}.$ (10)
In the limit of small victim load resistance ($R_{\text{v}}$), most of the energy is delivered to the spark, and $R_{\text{SF}}=R_{\text{SFmin}}$, so $\mathcal{E}_{\text{v}}=\frac{1}{2}\frac{R_{\text{v}}}{R_{\text{SFmin}}}\mathcal{E}_{0}$ and the fraction of stored energy delivered to the victim load is
$\eta_{\text{v}}=\frac{\mathcal{E}_{\text{v}}}{\mathcal{E}_{0}}=\frac{1}{2}\frac{R_{\text{v}}}{R_{\text{SFmin}}}=\frac{1}{4}a_{R}\,C\,R_{\text{v}}\,E_{\text{th}}^{2}.$ (11)
In a separate paper, we explore a more general solution for arbitrary $R_{\text{v}}$ [schrama2025variableRv]. Eq. (11) is the lowest-order term of that more general solution when expanded for small $R_{\text{v}}/R_{\text{SFmin}}$. When looking at the fractional energy, we can move all circuit parameters to one side to obtain
$\overline{\eta}_{\text{v}}=\frac{\eta_{\text{v}}}{CR_{\text{v}}}=\frac{a_{R}E_{\text{th}}^{2}}{4}.$ (12)
This expression shows that if the fraction of stored energy is scaled by $C\,R_{\text{v}}$, the result will be independent of the gap length. We will test this result in Section 4 below.